Hydrogen peroxide is a powerful oxidizing agent used in concentrated solution in rocket fuels and in dilutee solution as a hair bleach. An aqueous solution of H2O2is 30 % by mass and has a density of 1.11 g/mL. Calculate the Molality, Mole fraction of H2O2, and Molarity of the solution.

1
Expert's answer
2021-02-24T03:30:36-0500

Molality= m(solute)/m(solvent).

p=m/V.

Molality = 30/0.07= 428,571429.

30/34 = 0,882352941.

70/18 = 3,88888889 .

0.882352941 x 100/0.882352941+3,88888889 = 18,4931507 %. H2O2.

81,5068493% H2O.

Molarity=n/V=0.882352941 x 1.11/0.1=

9,79411765M.